OpenStudy (anonymous):

smith engineering found that an experienced surveyor surveys a roadbed in 4 hours. An apprentice surveyor needs 5 hours to survey the same stretch of road. If the two work together, find how long it takes them to complete the road.

OpenStudy (anonymous):

Turn it into an equation. So the experienced guy works at a pace of 1/4 per hour and the apprentice works at a pace of 1/5 per hour. The equation then becomes\[\frac{ 1 }{ 4 }x + \frac{ 1 }{ 5 }x = 1\] Where x is the hours worked and 1 is one whole road. Then solve the equation.

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would that equal 2/9x=1?

OpenStudy (anonymous):

I don't know if you mistyped but it's actually 9/20x = 1 To solve this, isolate the x by multiplying both side by 20/9.

OpenStudy (anonymous):

so x = 1*20/9

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 } + \frac{ 1 }{ 5 } = \frac{ 5 }{ 20} + \frac{ 4 }{ 20 } = \frac{ 9 }{ 20 }\]

OpenStudy (anonymous):

I'm confused on how that got to that.

OpenStudy (anonymous):

which part? the 9/20?

OpenStudy (anonymous):

the entire thing is just confusing to me. I apologize.

OpenStudy (anonymous):

No it's fine. Do you get how I got the original equation: 1/4x +1/5x = 1

OpenStudy (anonymous):

yes I understand that

OpenStudy (anonymous):

Okay now add the fractions 1/4x +1/5x. You can ignore the x for now and just add the fractions.

OpenStudy (anonymous):

1/20?

OpenStudy (anonymous):

When you add fractions, you have to change the fractions so that they have the same denominator (bottom number). To change the fractions, you have to multiply the top and bottom by the same or you change the value of the equation.In this case:\[\frac{ 1 }{ 4 }\times \frac{ 5 }{ 5 } + \frac{ 1 }{ 5 }\times \frac{ 4 }{ 4 }\]

OpenStudy (anonymous):

20 is the least common multiple of 4 and 5 (the smallest number that is a multiple of both 4 and 5) and you should change the denominator of both fractions to 20

OpenStudy (anonymous):

To change a fraction without changing its value, you need to multiply the top and bottom of the fraction by the same number

OpenStudy (anonymous):

Understand that part?

OpenStudy (anonymous):

yes I understand that

OpenStudy (anonymous):

Okay so now the fraction is 4/20 + 5/20 which is 9/20. Hence the 9/20x = 1

OpenStudy (anonymous):

Ok I got that part

OpenStudy (anonymous):

Now the solve for x, you need to get rid of the 9/20. The best way to do this is to multiply the fraction by its reciprocal (the fraction flipped upside down) In this case it would be 9/20 * 20/9 That would just leave x because of cross-multiplication. But remember, what you do to one side of an equation, you have to do to the other. So the equation become 9/20x * 20/9 = 1 * 20/9

OpenStudy (anonymous):

sorry not cross-multiplication. Ignore that part

OpenStudy (anonymous):

ok so you multiply 1* 20/9?

OpenStudy (anonymous):

yup which just 20/9

OpenStudy (anonymous):

so x = 20/9. It will take them both 20/9 hours to finish the road.

OpenStudy (anonymous):

oh alright well thank you for your help

OpenStudy (anonymous):

hope you learned something. it's hard to explain through typing

OpenStudy (anonymous):

I did thank you

OpenStudy (anonymous):

you're welcome